# Parabolic Equations with Singular Coefficients and Boundary Data: Analysis and Numerical Simulations

### v1

@@ 4. Very weak solutions
    $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\leq C|\log\varepsilon|^{N},\quad\|b_{\varepsilon}\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\leq C\varepsilon^{-N}.$
@@
The above assumptions are natural in the context of distributional coefficients, especially in one spatial dimension. Regularisations of distributions often form moderate nets. In particular, by the structure theorem for distributions (see, e.g., [16]), we have the inclusion
$\mathcal{D}^{\prime}(0,1)\subset\{\text{$L^{\infty}(0,1)$-moderate families}\}.$ (4.1)
This means that even if a solution to a boundary value problem on $(0,1)$ does not exist in the classical distributional sense, it may still exist as a net of $L^{\infty}$-moderate functions. This framework allows for a consistent treatment of singular coefficients and data.
@@
$\widetilde{f}(t,x):=\begin{cases}f(x),&(x)\in(0,1),\\ 0,&(x)\in\mathbb{R}\setminus(0,1).\end{cases}$
@@
  • Each $\widetilde{f}_{\varepsilon}\in C^{\infty}(\mathbb{R})$,
  • The net $(\widetilde{f}_{\varepsilon})$ is $L^{p}(\mathbb{R})$-moderate for any $1\leq p\leq\infty$,
  • On the original domain, $\widetilde{f}_{\varepsilon}\to f$ in $L^{p}(0,1)$ as $\varepsilon\to 0$,
@@
Definition 6.
@@
Assume that $a\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, $q\in\mathcal{D}^{\prime}(0,1)$, and $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$. A net of functions $(u_{\varepsilon})_{\varepsilon>0}$ is said to be a very weak solution to the initial/boundary value problem (4.2), if there exist an $L^{\infty}$-moderate regularisation $q_{\varepsilon}$ of $q$, a $W^{1,\infty}$-moderate regularisation $a_{\varepsilon}$ of $a$, a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation $b_{\varepsilon}$ of $b$, an $H^{1}$-moderate regularisation $u_{0,\varepsilon}$ of $u_{0}$, time-distributional moderate regularisations $g_{0,\varepsilon}$ and $g_{1,\varepsilon}$ of $g_{0}$ and $g_{1}$ respectively, and an $L^{2}$-moderate regularisation $f_{\varepsilon}$ of $f$, such that the regularised problem
@@
$\left\{\begin{array}[]{l}\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon})(t,x)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)=f_{\varepsilon}(t,x),\\ \hfill(t,x)\in[0,T]\times(0,1),\\ u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),\quad x\in(0,1),\\ u_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad u_{\varepsilon}(t,1)=g_{1,\varepsilon}(t),\quad t\in[0,T],\end{array}\right$. (4.3)
@@
holds for all $\varepsilon\in(0,1]$, and the net $(u_{\varepsilon})$ is moderate in the sense that there exist constants $C>0$, $N\in\mathbb{N}$ such that
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial^{2}_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\varepsilon^{-N},\quad\text{for all }\varepsilon\in(0,1].$
@@
Let the coefficients $q\in\mathcal{D}^{\prime}(0,1)$, $a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1))$, and the boundary data $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$. Then, the initial-boundary value problem (4.2) admits a very weak solution in the sense of Definition 6.

### v2

@@ 3. Very weak solutions for the non-homogeneous Dirichlet conditions
We additionally assume that the negative part $q_{-,\varepsilon}:=\max\{-q_{\varepsilon},0\}$ satisfies
$\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\leq C|\log\varepsilon|.$
@@
    $\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}\leq C|\log\varepsilon|,\quad\|b_{\varepsilon}\|_{L^{\infty}([0,T];W^{1,\infty}(0,1))}\leq C\varepsilon^{-N}.$
@@
The above moderateness assumptions are natural in the context of distributional coefficients and data, particularly in one spatial dimension. Let $v\in\mathcal{D}^{\prime}(0,1)$. After extension by zero to $\mathbb{R}$, the standard mollification $v_{\varepsilon}:=v*\psi_{\varepsilon}$ defines a net of smooth functions.
By the structure theorem for distributions (see, e.g., [19]), there exist constants $C>0$ and $N\in\mathbb{N}$, depending on the order of $v$, such that
$\|v_{\varepsilon}\|_{L^{\infty}(0,1)}\leq C\varepsilon^{-N},\qquad\varepsilon\in(0,1].$
Hence every distribution admits an $L^{\infty}$-moderate regularisation.
This observation justifies the use of moderate nets to represent distributional coefficients and data, even in situations where a classical distributional solution of the boundary value problem does not exist.
@@
$\widetilde{f}(x):=\begin{cases}f(x),&(x)\in(0,1),\\ 0,&(x)\in\mathbb{R}\setminus(0,1).\end{cases}$
@@
  • each $\widetilde{f}_{\varepsilon}\in C^{\infty}(\mathbb{R})$;
  • for each fixed $2\leq p\leq\infty$, the net $(\widetilde{f}_{\varepsilon})$ is $L^{p}(\mathbb{R})$-moderate;
  • on the original domain, $\widetilde{f}_{\varepsilon}\to f$ in $L^{p}(0,1)$ as $\varepsilon\to 0$;
@@
Definition 6 (Very weak solution).
Assume that
$a\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\qquad b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\qquad q\in\mathcal{D}^{\prime}(0,1),$
$f\in\mathcal{D}^{\prime}([0,T]\times(0,1)),\qquad u_{0}\in\mathcal{D}^{\prime}(0,1),\qquad g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T]).$
@@
A net of functions $(u_{\varepsilon})_{\varepsilon\in(0,1]}$ is said to be a very weak solution to the initial–boundary value problem (3.1) if there exist
  • an $L^{\infty}$-moderate regularisation $q_{\varepsilon}$ of $q$,
  • a $W^{1,\infty}$-moderate regularisation $a_{\varepsilon}$ of $a$,
  • a log-$L^{\infty}(W^{1,\infty})$-moderate regularisation $b_{\varepsilon}$ of $b$,
  • an $H^{1}$-moderate regularisation $u_{0,\varepsilon}$ of $u_{0}$,
  • time-distributional moderate regularisations $g_{0,\varepsilon}$, $g_{1,\varepsilon}$ of $g_{0}$, $g_{1}$, respectively,
  • an $L^{2}$-moderate regularisation $f_{\varepsilon}$ of $f$,
such that, for every $\varepsilon\in(0,1]$, the function $u_{\varepsilon}$ is the unique weak solution of the regularised problem
@@
$\left\{\begin{array}[]{l}\partial_{t}u_{\varepsilon}(t,x)-\partial_{x}\!\big(a_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)\big)+b_{\varepsilon}(t,x)\partial_{x}u_{\varepsilon}(t,x)+q_{\varepsilon}(x)u_{\varepsilon}(t,x)=f_{\varepsilon}(t,x),\\[2.0pt] \hfill(t,x)\in[0,T]\times(0,1),\\[2.0pt] u_{\varepsilon}(0,x)=u_{0,\varepsilon}(x),\qquad x\in(0,1),\\[2.0pt] u_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad u_{\varepsilon}(t,1)=g_{1,\varepsilon}(t),\qquad t\in[0,T],\end{array}\right$. (3.2)
where the regularised diffusion coefficient satisfies the uniform ellipticity condition
$a_{\varepsilon}(t,x)\geq\alpha>0,\qquad(t,x)\in[0,T]\times(0,1),\ \varepsilon\in(0,1].$
@@
Moreover, the net $(u_{\varepsilon})$ is required to be moderate in the sense that there exist constants $C>0$ and $N\in\mathbb{N}$ such that
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial_{x}^{2}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq C\,\varepsilon^{-N},\qquad\varepsilon\in(0,1].$
In the previous section, we considered the parabolic problem with regular coefficients and homogeneous Dirichlet boundary conditions. Here, we admit distributional coefficients and data, and, after regularisation, handle the non-homogeneous Dirichlet problem by homogenising the boundary conditions via a lifting.
We decompose the solution as
$u_{\varepsilon}(t,x)=w_{\varepsilon}(t,x)+\psi_{\varepsilon}(t,x),$ (3.3)
where $w_{\varepsilon}$ satisfies homogeneous Dirichlet boundary conditions
$w_{\varepsilon}(t,0)=w_{\varepsilon}(t,1)=0,$
and $\psi_{\varepsilon}$ is a lifting function defined by
$\psi_{\varepsilon}(t,x):=(1-x)g_{0,\varepsilon}(t)+xg_{1,\varepsilon}(t),\quad(t,x)\in[0,T]\times(0,1).$
This function satisfies:
  • $\psi_{\varepsilon}(t,0)=g_{0,\varepsilon}(t),\quad\psi_{\varepsilon}(t,1)=g_{1,\varepsilon}(t)$,
  • $\partial_{x}\psi_{\varepsilon}(t,x)=g_{1,\varepsilon}(t)-g_{0,\varepsilon}(t)$ is independent of $x$,
  • $\psi_{\varepsilon}\in H^{1}([0,T];H^{1}(0,1))$ provided $g_{0,\varepsilon},g_{1,\varepsilon}\in H^{1}([0,T])$.
Substituting (3.3) into (3.2), we obtain the equation for $w_{\varepsilon}$:
$\left\{\begin{array}[]{l}\partial_{t}w_{\varepsilon}-\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}w_{\varepsilon})+b_{\varepsilon}(t,x)\partial_{x}w_{\varepsilon}+q_{\varepsilon}(x)w_{\varepsilon}=\widetilde{f}_{\varepsilon}(t,x),\\[2.0pt] w_{\varepsilon}(0,x)=u_{0,\varepsilon}(x)-\psi_{\varepsilon}(0,x),\quad x\in(0,1),\\[2.0pt] w_{\varepsilon}(t,0)=w_{\varepsilon}(t,1)=0,\quad t\in[0,T],\end{array}\right$. (3.4)
where the inhomogeneous source term is given by
$\widetilde{f}_{\varepsilon}(t,x)=$ $f_{\varepsilon}(t,x)-\partial_{t}\psi_{\varepsilon}(t,x)+\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}\psi_{\varepsilon}(t,x))-b_{\varepsilon}(t,x)\partial_{x}\psi_{\varepsilon}(t,x)$
$-q_{\varepsilon}(x)\psi_{\varepsilon}(t,x)=f_{\varepsilon}(t,x)-(1-x)g_{0,\varepsilon}^{\prime}(t)-xg_{1,\varepsilon}^{\prime}(t)$
$+(\partial_{x}a_{\varepsilon}(t,x)-b_{\varepsilon}(t,x))(g_{1,\varepsilon}(t)-g_{0,\varepsilon}(t))-q_{\varepsilon}(x)\big((1-x)g_{0,\varepsilon}(t)+xg_{1,\varepsilon}(t)\big),$
where $\partial_{x}(a_{\varepsilon}(t,x)\partial_{x}\psi_{\varepsilon}(t,x))=(\partial_{x}a_{\varepsilon})(g_{1,\varepsilon}-g_{0,\varepsilon})$ since $\partial^{2}_{x}\psi_{\varepsilon}=0$.
The right-hand side $\widetilde{f}_{\varepsilon}\in L^{2}([0,T];L^{2}(0,1))$ under the assumptions
$f_{\varepsilon}\in L^{2}([0,T];L^{2}(0,1)),\quad a_{\varepsilon},\,\partial_{x}a_{\varepsilon},\,b_{\varepsilon}\in L^{\infty}([0,T];L^{\infty}(0,1))\quad q_{\varepsilon}\in L^{\infty}(0,1)$
and $g_{0,\varepsilon},g_{1,\varepsilon}\in H^{1}([0,T])$, and the fact that the coefficients are bounded. Moreover, using $H^{1}(0,T)\hookrightarrow L^{\infty}(0,T)$, we obtain
$\|\widetilde{f}_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}\leq\;$ $\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|g^{\prime}_{0,\varepsilon}\|_{L^{2}(0,T)}+\|g^{\prime}_{1,\varepsilon}\|_{L^{2}(0,T)}$
$+\Big(\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T]\times(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T]\times(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\Big)$
$\qquad\times\Big(\|g_{0,\varepsilon}\|_{L^{\infty}(0,T)}+\|g_{1,\varepsilon}\|_{L^{\infty}(0,T)}\Big)$
$\leq\;$ $\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+C\Big(1+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}}+\|b_{\varepsilon}\|_{L^{\infty}}+\|q_{\varepsilon}\|_{L^{\infty}}\Big)$
$\qquad\times\Big(\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\Big),$ (3.5)
where the constant $C$ depends only on $T$.
Estimate of the initial data.
Since
$\psi_{\varepsilon}(0,x)=(1-x)g_{0,\varepsilon}(0)+xg_{1,\varepsilon}(0),$
we have by the triangle inequality
$\|w_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}=\|u_{0,\varepsilon}-\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}\leq\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}.$
Moreover,
$\|\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}^{2}=\int_{0}^{1}\big|(1-x)g_{0,\varepsilon}(0)+xg_{1,\varepsilon}(0)\big|^{2}\,dx\leq C\big(|g_{0,\varepsilon}(0)|^{2}+|g_{1,\varepsilon}(0)|^{2}\big),$
hence
$\|\psi_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}\leq C\big(|g_{0,\varepsilon}(0)|+|g_{1,\varepsilon}(0)|\big).$
Finally, by the trace inequality $|h(0)|\leq C\|h\|_{H^{1}(0,T)}$ (valid for all $h\in H^{1}(0,T)$), we obtain
$\|w_{\varepsilon}(0,\cdot)\|_{L^{2}(0,1)}\leq\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+C\big(\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\big),$ (3.6)
where $C>0$ depends only on $T$ and the interval $(0,1)$.
Writing $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, the function $w_{\varepsilon}$ is the unique weak solution of the homogeneous problem (3.4) with initial datum $u_{0,\varepsilon}-\psi_{\varepsilon}(0)$ and source term $\widetilde{f}_{\varepsilon}$.
Proposition 3.1.
Let $u_{\varepsilon}$ be the weak solution to (3.2) with $u_{0,\varepsilon}\in L^{2}(0,1)$ and $f_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$. Assume
$a_{\varepsilon}(t,x)\geq\alpha>0,\qquad a_{\varepsilon}\in L^{\infty}(0,T;L^{\infty}(0,1)),\qquad b_{\varepsilon}\in L^{\infty}(0,T;W^{1,\infty}(0,1)),$
and
$q_{\varepsilon}\in L^{\infty}(0,1),\qquad q_{\varepsilon}=q_{+,\varepsilon}-q_{-,\varepsilon},\quad q_{+,\varepsilon}\geq 0,\quad q_{-,\varepsilon}:=\max\{-q_{\varepsilon},0\}.$
Then $u_{\varepsilon}\in L^{2}(0,T;H^{1}_{0}(0,1))$ and $\partial_{t}u_{\varepsilon}\in L^{2}(0,T;H^{-1}(0,1))$, and
$\|u_{\varepsilon}\|_{L^{\infty}(0,T;L^{2}(0,1))}+\sqrt{\alpha}\,\|\partial_{x}u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}\,u_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}$
$\qquad\leq C\Big(\exp\!\left\{T\big(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\big)\right\}$
$\qquad\quad\times\big(1+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\big)$
$\qquad\quad\times\Big(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\Big)$
$\qquad\quad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big),$ (3.7)
where the constant $C$ depends only on $T$ and the domain $(0,1)$, but not on the data or coefficients.
The estimate (3.1) shows that $u_{\varepsilon}$ is a moderate net whenever the regularisations of the coefficients and data are moderate in the sense of Definition 6.
Proof.
We write $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, where $w_{\varepsilon}$ satisfies the homogeneous problem (3.4). Applying the energy estimate from Theorem 2.2, we obtain
$\|w_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad\leq C\exp{\left\{T\big(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\big)\right\}}$
$\qquad\quad\times\big(\|u_{0,\varepsilon}-\psi_{\varepsilon}(0)\|_{L^{2}(0,1)}+\|\widetilde{f}_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\big).$
Taking into account (3) and (3.6), we obtain
$\|w_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad\leq C\exp{\left\{T\big(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\big)\right\}}$
$\qquad\quad\times\Big(1+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}\Big)$
$\qquad\quad\times\big(\|u_{0,\varepsilon}\|_{L^{2}(0,1)}+\|f_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}(0,T)}+\|g_{1,\varepsilon}\|_{H^{1}(0,T)}\big).$ (3.8)
Next, since $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, we use the triangle inequality as follows:
$\|u_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}\leq\|w_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\|\psi_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))},$
$\sqrt{\alpha}\|\partial_{x}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq\sqrt{\alpha}\|\partial_{x}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))},$
$\|\sqrt{q_{+,\varepsilon}}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}\leq\|\sqrt{q_{+,\varepsilon}}w_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}.$
Note that $\psi_{\varepsilon}(t,x)\in\operatorname{span}\{1,x\}\cdot g_{j,\varepsilon}(t)$, so
$\|\psi_{\varepsilon}\|_{L^{\infty}([0,T];L^{2}(0,1))}+\sqrt{\alpha}\|\partial_{x}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\sqrt{q_{+,\varepsilon}}\psi_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\qquad\leq C(T)\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\big).$
Combining this with (3) yields (3.1), which completes the proof of Proposition 3.1. ∎
Remark 3.2.
The last term in (3.1) corresponds to the explicit contribution of the lifting function $\psi_{\varepsilon}$; since $\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}$ is $L^{\infty}$-moderate and $g_{0,\varepsilon},g_{1,\varepsilon}$ are $H^{1}$-moderate, this contribution is moderate as well.
Proposition 3.2.
Let $u_{\varepsilon}$ solve the regularized non-homogeneous Dirichlet problem (3.2) with data
$f_{\varepsilon}\in L^{2}([0,T];L^{2}(0,1)),\quad u_{0,\varepsilon}\in H^{1}_{0}(0,1),\quad g_{0,\varepsilon},\,g_{1,\varepsilon}\in H^{1}([0,T]).$
Assume the coefficients satisfy
$a_{\varepsilon},\,\partial_{t}a_{\varepsilon},\,\partial_{x}a_{\varepsilon},\,b_{\varepsilon},\,\partial_{x}b_{\varepsilon}\in L^{\infty}([0,T];L^{\infty}(0,1)),\quad a_{\varepsilon}(t,x)\geq\alpha>0,$
$q_{\varepsilon}\in L^{\infty}(0,1),\qquad q_{\varepsilon}=q_{+,\varepsilon}-q_{-,\varepsilon},\quad q_{+,\varepsilon}\geq 0,\quad q_{-,\varepsilon}:=\max\{-q_{\varepsilon},0\}.$
Then $\partial_{t}u_{\varepsilon},\,\partial_{x}(a_{\varepsilon}\partial_{x}u_{\varepsilon})\in L^{2}([0,T];L^{2}(0,1))$, and the estimate
$\|\partial_{t}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial_{x}(a_{\varepsilon}\partial_{x}u_{\varepsilon})\|_{L^{2}([0,T];L^{2}(0,1))}+\|\partial_{x}^{2}u_{\varepsilon}\|_{L^{2}([0,T];L^{2}(0,1))}$
$\quad\leq\frac{C}{\alpha}\Big(\exp{\left\{T\left(\|\partial_{x}b_{\varepsilon}\|_{L^{\infty}([0,T];L^{\infty}(0,1))}+2\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\right)\right\}}\Big(1+\|q_{\varepsilon}\|_{L^{\infty}(0,1)}$
$\qquad+\frac{1}{\sqrt{\alpha}}\big(\|\partial_{t}a_{\varepsilon}\|^{1/2}_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|\partial_{x}a_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}+\|b_{\varepsilon}\|_{L^{\infty}(0,T;L^{\infty}(0,1))}\big)\Big)$
$\qquad\times\Big(\sqrt{\alpha}\|u_{0,\varepsilon}\|_{H^{1}_{0}(0,1)}+\|f_{\varepsilon}\|_{L^{2}(0,T;L^{2}(0,1))}+\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\Big)$
$\qquad+\big(1+\sqrt{\alpha}+\|\sqrt{q_{+,\varepsilon}}\|_{L^{\infty}(0,1)}\big)\big(\|g_{0,\varepsilon}\|_{H^{1}([0,T])}+\|g_{1,\varepsilon}\|_{H^{1}([0,T])}\big)\Big),$ (3.9)
where $C>0$ depends only on $T$ and the interval $(0,1)$.
Proof.
We decompose $u_{\varepsilon}=w_{\varepsilon}+\psi_{\varepsilon}$, where $\psi_{\varepsilon}$ is the boundary lifting and $w_{\varepsilon}$ satisfies homogeneous Dirichlet conditions.
Then $w_{\varepsilon}$ solves the homogeneous problem (3.4) with source term
$\widetilde{f}_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$
and initial datum $u_{0,\varepsilon}-\psi_{\varepsilon}(0)\in H^{1}_{0}(0,1)$. Hence Lemma 2.4 applies and yields
$\partial_{t}w_{\varepsilon},\ \partial_{x}(a_{\varepsilon}\partial_{x}w_{\varepsilon})\in L^{2}(0,T;L^{2}(0,1)).$
Using the triangle inequality and the explicit form of $\psi_{\varepsilon}$, we obtain the corresponding bounds for $u_{\varepsilon}$. Moreover, since $\partial_{x}u_{\varepsilon}=\partial_{x}w_{\varepsilon}+\partial_{x}\psi_{\varepsilon}$ and $\partial_{x}\psi_{\varepsilon}\in L^{2}(0,T)$, we have $\partial_{x}u_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$.
Finally,
$a_{\varepsilon}\partial_{x}^{2}u_{\varepsilon}=\partial_{x}(a_{\varepsilon}\partial_{x}u_{\varepsilon})-(\partial_{x}a_{\varepsilon})\partial_{x}u_{\varepsilon},$
and $a_{\varepsilon}\geq\alpha>0$, $\partial_{x}a_{\varepsilon}\in L^{\infty}$ imply $\partial_{x}^{2}u_{\varepsilon}\in L^{2}(0,T;L^{2}(0,1))$. This yields (3.2). ∎
@@
Let
$a,\,b\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\quad q\in\mathcal{D}^{\prime}(0,1),\quad f\in\mathcal{D}^{\prime}([0,T];\mathcal{D}^{\prime}(0,1)),\quad u_{0}\in\mathcal{D}^{\prime}(0,1),$
and let the boundary data satisfy $g_{0},g_{1}\in\mathcal{D}^{\prime}([0,T])$. Assume that there exist regularisations
$(a_{\varepsilon}),\ (b_{\varepsilon}),\ (q_{\varepsilon}),\ (f_{\varepsilon}),\ (u_{0,\varepsilon}),\ (g_{0,\varepsilon}),\ (g_{1,\varepsilon})$
as in Definition 6, with $a_{\varepsilon}(t,x)\geq\alpha>0$ uniformly in $\varepsilon$, and such that $q_{\varepsilon}$ is log-$L^{\infty}$-moderate in the sense that
$\|q_{-,\varepsilon}\|_{L^{\infty}(0,1)}\leq C|\log\varepsilon|.$
Then the problem (3.1) admits a very weak solution in the sense of Definition 6.